OpenStudy (anonymous):

f(x)=(ax+b)/(cx+d), abcd are constants. If ad=bc, show that f is a constant function.

OpenStudy (anonymous):

oh here

OpenStudy (anonymous):

we can do this quickly if we get to use calculus

OpenStudy (anonymous):

yes we can

OpenStudy (anonymous):

ok then take the derivative, say "oh look it is zero" and therefore your function must be a constant. would you like me to write it out?

OpenStudy (anonymous):

if you have time thx

OpenStudy (anonymous):

\[f'(x)=\frac{(cx+d)a-(ax+b)c}{(cx+c)^2}\] \[=\frac{acx+ad-acx-bc}{(cx+d)^2}=\frac{ad-bc}{(cx+d)^2}\] and since \[ad-bc=0\] you get \[f'(x)=0\] for all x making f a constant

OpenStudy (anonymous):

k thx dude